### v3

@@ 4.2 Mid-Level Motion Planning
In this framework, the LLMs are limited to allocating unexplored frontiers to each robot for navigation, serving primarily as a task allocation mechanism. Morad et al. [68] took a step further and proposed a novel framework combining LLMs with offline reinforcement learning (RL) to address path-finding challenges in MRS. Their approach involves leveraging LLMs to translate natural language commands into latent embeddings, which are then encoded with agent observations to create state-task representations. Using offline RL, policies are trained on these representations to generate navigation strategies that understand and follow high-level natural language tasks. A key advantage of this framework is its ability to train policies entirely on real-world data without requiring simulators, ensuring direct applicability to physical robots. The integration of LLMs enhances the flexibility of task instruction interpretation, while RL facilitates the generation of low-latency and reactive control policies, enabling efficient multi-robot navigation. Following this line, Godfrey et al. [26] developed MARLIN (Multi-Agent Reinforcement Learning Guided by Language-Based Inter-Robot Negotiation), a framework combining LLMs with Multi-Agent Proximal Policy Optimization (MAPPO) to enhance training efficiency and transparency in multi-robot navigation tasks. In MARLIN, robots equipped with LLMs engage in natural language negotiations to collaboratively generate task plans, which are then used to guide policy training. This hybrid approach dynamically switches between LLM-guided planning and standard MAPPO-based reinforcement learning, leveraging the reasoning capabilities of LLMs to improve training speed and sample efficiency without sacrificing performance. Experimental results demonstrate MARLIN’s ability to achieve faster convergence and more consistent performance compared to conventional MARL approaches, with applications validated in both simulation and physical robot environments. This integration of negotiation-based planning highlights the potential of combining LLMs with MARL for scalable, explainable multi-robot coordination. On the other hand, Garg et al. [24] leveraged LLMs for deadlock resolution in connected multi-robot navigation systems. In obstacle-laden environments, such systems can experience deadlocks that low-level control policies cannot resolve. To address this, the LLM selects a leader robot and plans waypoints for it to reach the target. The system reconfigures into a leader-follower formation, with a GNN-based low-level controller guiding the leader along the waypoints. Similarly, Wu et al. [97] proposed a mid-level action LLM that uses short-horizon inputs, such as tracking errors and control costs, to generate parameters for an optimization-based robot controller, enabling it to follow planned waypoints effectively. While the aforementioned research primarily employs centralized systems where LLMs handle planning for all robots, Wu et al. [96] developed a decentralized multi-robot navigation system for household tasks. In this framework, each robot is equipped with an LLM to enable communication and collaboration. Robots dynamically recognize and approach target objects distributed across multiple rooms. Leadership is dynamically assigned through a communication-triggered mechanism, with the leader robot issuing orders based on the global information it gathers. This flexible and decentralized leadership strategy enhances adaptability and efficiency in collaborative navigation scenarios.
@@ 4.3 Low-Level Action Generation
Strobel et al. [83] introduced LLM2Swarm, a system that integrates LLMs with robot swarms through two approaches: centralized controller synthesis and decentralized direct integration. In the centralized approach, LLMs are used to design and validate controllers prior to deployment, enabling efficient and adaptive behavior generation. In the decentralized approach, each robot has its own LLM instance, enabling localized reasoning, planning, and collaboration for enhanced flexibility in dynamic environments. The results highlight the potential of LLMs in swarm robotics, demonstrating their applicability to both centralized and decentralized control paradigms. Lykov et al. [63] further showcased the potential of LLMs in swarm control with FlockGPT, a framework for orchestrating UAV flocks to achieve desired geometric formations. In this system, the LLM generates a signed distance function (SDF) to guide UAV movements relative to a target surface, while a dedicated control algorithm manages practical constraints such as collision avoidance. These studies underscore the versatility of LLMs in enhancing both centralized and decentralized swarm behaviors.
@@ 4.4 Human Intervention
In MRS, LLMs typically focus on executing tasks based on human-provided instructions, emphasizing the interpretation of instructions and autonomous task completion. Once the instructions are delivered, human involvement is often minimized. However, emerging research explores scenarios that require continuous interaction between LLMs and humans, emphasizing cooperation, decision-making, or external observation throughout task execution. These studies highlight the potential for dynamic human intervention to address unexpected challenges, refine task strategies, or ensure safety in critical applications. By enabling iterative human-robot collaboration, such approaches enhance the adaptability and reliability of LLM-driven MRS. The simplest form of human-robot interaction is demonstrated by Lakhnati et al. [45], where robots operate in a straightforward cycle: receiving a human command, executing the corresponding task, reporting the completion status, and awaiting the next instruction. Building on this, Lykov et al. [62] introduced the LLM-MARS framework, which enables humans to inquire about each robot’s current state and task progress at any time. In this system, both response generation and task execution are handled by a single LLM, enhanced with distinct LoRA adapters for efficiency. Hunt et al. [36] proposed a more interactive approach, requiring human approval before executing any plan generated through LLM-driven discussions. If the proposed plan is deemed unreasonable, the human supervisor can provide feedback, prompting the LLMs to refine their approach through further dialogue. Ahn et al. [2] introduced the VADER system, further enhancing human involvement. When a robot encounters a task-related issue, it posts a request for assistance on the Human-Robot Fleet Orchestration Service (HRFS), a shared platform accessible to both human operators and robotic agents. Any agent or human can respond to the request, and once the issue is resolved, the robot resumes its task. These examples illustrate the varying degrees of human involvement in LLM-driven MRS, ranging from simple command execution to active collaboration and dynamic problem-solving.
@@
High level Task Allocation and Task Planning Wu et al. [97] Cent Homo T Sim(ROS,RViz), Real Target Tracking GPT-4o, GPT-3.5 Turbo Brienza et al. [9] Cent Homo T, I Sim(SimRobot) Game V: GPT-4 Turbo, L: GPT-3.5 Turbo Lykov et al. [62] Cent Homo T, A Real General Purpose Falcon Ahn et al. [2] Dec Hetero T, I Real General Purpose V: CLIP, ViLD, PaLI, L: PaLM Lakhnati et al. [45] Hier Hetero T, A Sim(VR) General Purpose GPT-4 Chen et al. [12] Hier Hetero T Sim Household GPT-4o Chen et al. [15] Hier Hetero T Sim(AI2THOR) General Purpose GPT-4, GPT-3.5 Turbo Gupte et al. [31] Cent Hetero T Sim(GAMA) General Purpose - Huang et al. [35] Cent Homo T Sim General Purpose GPT-4 Turbo Kannan et al. [40] Cent Homo T Sim(AI2THOR), Real Household GPT-4, GPT-3.5, Llama2, Claude3 Wang et al. [94] Cent Hetero T, I Sim Construction L: Llama-3.1, Claude 3.5, GPT-4o,GPT-3.5 Turbo, V: CLIP Xu et al. [99] Cent Homo T Sim(AI2THOR+ALFRED), Real Household Minstral, GPT-4 Obata et al. [70] Cent Hetero T Sim(ROS) General Purpose GPT Liu et al. [57] Hier Hetero T Sim(BEHAVIOR-1K), Real General Purpose GPT-4 Mandi et al. [65] Dec Homo T Sim,Real Household GPT-4 Yu et al. [105] Dec Hetero T Sim(PyBullet) Household GPT-3.5 Turbo, GPT-4o, Llama3.1 Sueoka et al. [84] Dec Hetero T, I Real Construction V: GPT-4v, L: GPT-4 Hunt et al. [36] Cent Homo T Real General Purpose GPT-4 Yoshida et al. [103] Dec Hetero T, I - General Purpose - Wang et al. [93] Cent Hetero T Sim(AI2THOR) General Purpose GPT-3.5, Llama-2, Llama-3 Guzman-Merino et al. [32] Cent Hetero T - General Purpose GPT-4o Mid level Motion Planning Yu et al. [104] Cent Homo T Sim(HM3D) Household GPT-3.5 Turbo Morad et al. [68] Dec Homo T Real General Purpose GTE-Based Godfrey et al. [26] Cent Homo T Sim, Real General Purpose Llama-3.1 Garg et al. [24] Cent Homo T, I Sim, Real General Purpose V: Claude-3 Sonnet/Opus, GPT-4 Turbo, GPT-4o, L: GPT-4, GPT-3.5, Claude-2, Claude-3 Wu et al. [97] Cent Homo T Sim(ROS, RViz), Real Target Tracking GPT-4o, GPT-3.5 Turbo Wu et al. [96] Dec Homo T Sim Household GPT-4o Low level Action Execution Chen et al. [14] Cent Homo T Sim General Purpose GPT-4 Venkatesh et al. [89] Cent Homo T, I Sim(Pygame), Real General Purpose GPT-4, Llama-2, Claude-3 Opus Li et al. [49] Dec Homo T Sim General Purpose GPT-3.5 Turbo Strobel et al. [83] Dec Homo T Sim(ARGoS), Real General Purpose GPT-3.5, GPT-4 Lykov et al. [63] Cent Homo T Sim Formation GPT-4
@@ 5 Application
Others including Construction, Formation, Target Tracking and Game. Several studies focused on applications in open-world environments, highlighting the versatility and innovative potential of LLM-integrated robotic systems. For instance, Wang et al. [94] and Sueoka et al. [84] explored the use of LLMs in orchestrating robotic systems for excavation and transportation tasks, showcasing their applicability in construction and complex terrain rescue operations. In drone formations applications, Lykov et al. [63] emphasized the coordination and adaptability needed for outdoor tasks such as search-and-rescue missions and environmental monitoring. Similarly, Wu et al. [97] addressed open-world target tracking by integrating danger zone recognition, providing robust solutions for autonomous navigation in hazardous environments. These scenarios further demonstrate the potential of LLMs in dynamic and structured environments. Brienza et al. [9] introduced LLCoach, a framework for robotic soccer applications, where LLMs enhance strategic decision-making and team coordination. Collectively, these studies underscore the potential of LLM-driven MRS to tackle diverse and complex challenges across various domains.
@@ 6 LLMs, Simulations, and Benchmarks
6.1 LLMs and VLMs
LLMs and VLMs play an increasingly significant role in MRS by enabling advanced decision-making, communication, and perception-driven collaboration. Different models offer unique strengths, making them suitable for specific MRS applications. Table 1 provides a comparative summary of LLMs and VLMs used in the studies discussed earlier, highlighting their contributions to multi-robot coordination, planning, and perception. GPT is one of the most extensively used language models, as demonstrated in Table 1, where it forms the backbone of many referenced studies. Its general-purpose reasoning and adaptability allow it to be integrated into multi-robot coordination tasks such as task allocation and planning, multi-robot communication, and human-robot collaboration [97, 35, 57, 45, 36]. Furthermore, GPT has been extended to VLM for applications requiring the integration of textual and visual inputs [24, 9, 84]. By incorporating fine-tuning techniques and vision encoders, GPT can analyze images, generate detailed descriptions, and seamlessly combine textual reasoning with visual comprehension. These capabilities make it well-suited for tasks such as image captioning, visual question answering, and multi-modal translation, expanding its utility into complex, perception-driven applications [9, 84, 24, 89]. Llama offers a range of open-sourced models from lightweight smaller models to powerful large-scale ones, catering to diverse application needs. The smaller models, such as Llama 3-8B, are particularly favored for their lightweight design and flexibility, making them highly suitable for embedded or decentralized MRS architectures where computational resources are limited. On the other hand, larger models like Llama 3-70B provide enhanced capabilities and higher accuracy, making them well-suited for complex tasks requiring advanced reasoning and detailed natural language understanding. Claude, on the other hand, prioritizes safety, ethical AI, and transparent decision-making, making it well-suited for regulated multi-robot applications. Additionally, Claude has been extended to function as a VLM [24], further expanding its versatility. Its strong focus on safety and ethical considerations makes Claude a compelling choice for tasks involving sensitive visual data, such as medical imaging or content moderation. Claude’s VLM implementations adopt a human-centric design, emphasizing transparency in decision-making and minimizing biases in visual interpretations. Unlike GPT, which often employs modular designs for vision-language tasks, Claude integrates safety protocols deeply into its VLM capabilities, making it well-suited for regulated industries requiring robust guardrails. GPT, Llama, and Claude exhibit a degree of interchangeability, as evidenced by several studies [40, 93, 89, 94] that tested architectures with multiple models, enabling comparative analysis of their performance. Falcon, with its emphasis on practicality, is optimized for resource-constrained environments. For instance, the only study utilizing Falcon [62] selected it as the preferred model due to computational limitations on each robot, which operated with microcomputers. PaLM stands out for its multi-tasking and multi-modal capabilities, excelling in complex reasoning and cross-domain tasks such as translation and image processing. However, its partially closed-source nature and integration within Google’s ecosystem make it predominantly utilized in Google DeepMind’s research [2]. Moreover, several VLMs have been explored in recent studies [2], including PaLI, CLIP, and ViLD. PaLI, developed by Google, is a multi-modal model tailored for multilingual and cross-visual tasks such as image captioning and visual question answering, leveraging extensive multi-modal data. Similarly, CLIP, created by OpenAI, aligns images and text within a shared embedding space through contrastive learning, making it particularly effective for zero-shot tasks like image classification and retrieval. In contrast, ViLD, another model from Google, specializes in zero-shot object detection by integrating vision features with CLIP-style language alignment, enabling it to precisely identify unseen object categories.

### v4

@@ 4.2 Mid-Level Motion Planning
In this framework, the LLMs are limited to allocating unexplored frontiers to each robot for navigation, serving primarily as a task allocation mechanism. Morad et al. [88] took a step further and proposed a novel framework combining LLMs with offline reinforcement learning (RL) to address path-finding challenges in MRS. Their approach involves leveraging LLMs to translate natural language commands into latent embeddings, which are then encoded with agent observations to create state-task representations. Using offline RL, policies are trained on these representations to generate navigation strategies that understand and follow high-level natural language tasks. A key advantage of this framework is its ability to train policies entirely on real-world data without requiring simulators, ensuring direct applicability to physical robots. The integration of LLMs enhances the flexibility of task instruction interpretation, while RL facilitates the generation of low-latency and reactive control policies, enabling efficient multi-robot navigation. Following this line, Godfrey et al. [31] developed MARLIN (Multi-Agent Reinforcement Learning Guided by Language-Based Inter-Robot Negotiation), a framework combining LLMs with Multi-Agent Proximal Policy Optimization (MAPPO) to enhance training efficiency and transparency in multi-robot navigation tasks. In MARLIN, robots equipped with LLMs engage in natural language negotiations to collaboratively generate task plans, which are then used to guide policy training. This hybrid approach dynamically switches between LLM-guided planning and standard MAPPO-based reinforcement learning, leveraging the reasoning capabilities of LLMs to improve training speed and sample efficiency without sacrificing performance. Experimental results demonstrate MARLIN’s ability to achieve faster convergence and more consistent performance compared to conventional MARL approaches, with applications validated in both simulation and physical robot environments. This integration of negotiation-based planning highlights the potential of combining LLMs with MARL for scalable, explainable multi-robot coordination. On the other hand, Garg et al. [29] leveraged LLMs for deadlock resolution in connected multi-robot navigation systems. In obstacle-laden environments, such systems can experience deadlocks that low-level control policies cannot resolve. To address this, the LLM selects a leader robot and plans waypoints for it to reach the target. The system reconfigures into a leader-follower formation, with a GNN-based low-level controller guiding the leader along the waypoints. Mahadevan et al. [83] directly tackled the challenge of multi-robot deadlocks by introducing GameChat, a role-based language negotiation framework for collaborative decision-making. In this system, agents engage in natural language dialogues to dynamically assign roles—such as “leader” or “blocker”—which help resolve spatial conflicts, task interdependencies, and bottlenecks that typically lead to deadlocks. These roles emerge through decentralized negotiation, without requiring pre-defined hierarchies or centralized control. The framework demonstrates robust performance across a variety of collaborative tasks, including navigation and manipulation, in both simulation and real-world experiments. GameChat showcases how emergent role-based reasoning and interactive dialogue can effectively address deadlock resolution in complex multi-agent environments. Moreover, Wu et al. [124] proposed a mid-level action LLM that uses short-horizon inputs, such as tracking errors and control costs, to generate parameters for an optimization-based robot controller, enabling it to follow planned waypoints effectively. While the aforementioned research primarily employs centralized systems where LLMs handle planning for all robots, Wu et al. [123] developed a decentralized multi-robot navigation system for household tasks. In this framework, each robot is equipped with an LLM to enable communication and collaboration. Robots dynamically recognize and approach target objects distributed across multiple rooms. Leadership is dynamically assigned through a communication-triggered mechanism, with the leader robot issuing orders based on the global information it gathers. This flexible and decentralized leadership strategy enhances adaptability and efficiency in collaborative navigation scenarios. Shen et al. [104] proposed a framework of decentralized semantic planning based on multimodal chain-of-thought reasoning. The MCoCoNav architecture consists of three modules and two auxiliary components. The Perception module evaluates the exploration value of the current scene using multimodal CoT reasoning; the Judgment module determines whether a robot should explore a new frontier or revisit a history node by analyzing the global semantic map; the Decision module selects the next long-term navigation goal based on predicted frontier scores or historical node scores. If the robots are stuck or trapped in place, the system resets their goal to a randomly sampled point on the global map to avoid deadlocks and ensure continued exploration. Ji et al. [46] proposed a reinforcement learning framework that groups LLM planners with collision and reachability constraints, ensuring physically valid multi-robot motion plans. Evaluations in two BoxNet environments demonstrate that constraint-aware small-scale LLMs significantly outperform larger ungrounded models in task success and generalization, highlighting the effectiveness of explicit constraint integration for reliable multi-robot planning. Rajvanshi et al. [96] introduced SayCoNav, a LLM-based decentralized adaptive planning approach. By dynamically updating collaboration strategies among heterogeneous robots, SayCoNav significantly improves multi-object search efficiency and adapts to changing robot conditions during mission execution, as validated through extensive simulation studies. Wang et al. [119] presented SAMALM, a decentralized multi-agent LLM actor-critic framework that ensures socially aware robot navigation. Utilizing parallel LLM actors and critics for verification and cooperative navigation, the system effectively balances local autonomy and global coordination, validated through diverse multi-robot navigation scenarios.
@@ 4.3 Low-Level Action Generation
In the field of formation control and planning, Strobel et al. [107] introduced LLM2Swarm, a system that integrates LLMs with robot swarms through two approaches: centralized controller synthesis and decentralized direct integration. In the centralized approach, LLMs are used to design and validate controllers prior to deployment, enabling efficient and adaptive behavior generation. In the decentralized approach, each robot has its own LLM instance, enabling localized reasoning, planning, and collaboration for enhanced flexibility in dynamic environments. The results highlight the potential of LLMs in swarm robotics, demonstrating their applicability to both centralized and decentralized control paradigms. Lykov et al. [80] further showcased the potential of LLMs in swarm control with FlockGPT, a framework for orchestrating UAV flocks to achieve desired geometric formations. In this system, the LLM generates a signed distance function (SDF) to guide UAV movements relative to a target surface, while a dedicated control algorithm manages practical constraints such as collision avoidance. These studies underscore the versatility of LLMs in enhancing both centralized and decentralized swarm behaviors. Moreover, Xue et al. [129] proposed a novel framework that leverages LLMs to achieve communication-efficient multi-robot formation control. The approach encodes formation goals, robot roles, and spatial arrangements into compact natural language instructions, which are then distributed to individual robots. These instructions are decoded into actionable local goals by each agent. The framework demonstrates that LLM-generated descriptions can maintain formation integrity while significantly lowering communication bandwidth. Following this line, Ji et al. [47] proposed GenSwarm, an end-to-end framework that enables automatic code-policy generation and deployment for multi-robot systems. The system interprets natural language instructions to construct constraint-based skill graphs, then hierarchically generates executable code policies using LLM agents. These policies are automatically reviewed, tested in simulation, and deployed to real-world robots via scalable hardware and software infrastructure. By supporting zero-shot generalization and human/VLM-in-the-loop feedback, GenSwarm offers a highly adaptive and interpretable solution for dynamic multi-robot tasks. Moreover, it is worth mentioning that GenSwarm supports optional human intervention in the loop, allowing users to provide natural language feedback after observing execution outcomes. This feedback is interpreted by LLM agents to iteratively refine control policies, enabling intuitive human-in-the-loop policy optimization.
@@ 4.4 Human Intervention
In MRS, LLMs typically focus on executing tasks based on human-provided instructions, emphasizing the interpretation of instructions and autonomous task completion. Once the instructions are delivered, human involvement is often minimized. However, emerging research explores scenarios that require continuous interaction between LLMs and humans, emphasizing cooperation, decision-making, or external observation throughout task execution. These studies highlight the potential for dynamic human intervention to address unexpected challenges, refine task strategies, or ensure safety in critical applications. By enabling iterative human-robot collaboration, such approaches enhance the adaptability and reliability of LLM-driven MRS. The simplest form of human-robot interaction is demonstrated by Lakhnati et al. [56], where robots operate in a straightforward cycle: receiving a human command, executing the corresponding task, reporting the completion status, and awaiting the next instruction. Building on this, Lykov et al. [79] introduced the LLM-MARS framework, which enables humans to inquire about each robot’s current state and task progress at any time. In this system, both response generation and task execution are handled by a single LLM, enhanced with distinct LoRA adapters for efficiency. Hunt et al. [43] proposed a more interactive approach, requiring human approval before executing any plan generated through LLM-driven discussions. If the proposed plan is deemed unreasonable, the human supervisor can provide feedback, prompting the LLMs to refine their approach through further dialogue. Ahn et al. [2] introduced the VADER system, further enhancing human involvement. When a robot encounters a task-related issue, it posts a request for assistance on the Human-Robot Fleet Orchestration Service (HRFS), a shared platform accessible to both human operators and robotic agents. Any agent or human can respond to the request, and once the issue is resolved, the robot resumes its task. Li et al. [66] introduced HMCF, which integrates LLM-driven multi-robot coordination with human oversight. Their framework mitigates hallucination via human verification, significantly improving task success rates and providing robust zero-shot generalization capabilities in diverse robotic tasks. These examples illustrate the varying degrees of human involvement in LLM-driven MRS, ranging from simple command execution to active collaboration and dynamic problem-solving.
@@
High level Task Allocation and Task Planning
Wu et al. [124] Cent Homo T Sim(ROS,RViz), Real Target Tracking GPT-4o, GPT-3.5 Turbo
Brienza et al. [10] Cent Homo T, I Sim(SimRobot) Game V: GPT-4 Turbo, L: GPT-3.5 Turbo
Lykov et al. [79] Cent Homo T, A Real General Purpose Falcon
Ahn et al. [2] Dec Hetero T, I Real General Purpose V: CLIP, ViLD, PaLI, L: PaLM
Lakhnati et al. [56] Hier Hetero T, A Sim(VR) General Purpose GPT-4
Chen et al. [15] Hier Hetero T Sim Household GPT-4o
Lim et al. [69] Cent Homo T, I Real Manufacturing GPT-4o
Jiang et al. [48] Dec Homo T Sim General Purpose GPT-4o, Gemini-2.0-Flash, DeepSeek-V3
Chen et al. [18] Hier Hetero T Sim(AI2THOR) General Purpose GPT-4, GPT-3.5 Turbo
Gupte et al. [37] Cent Hetero T Sim(GAMA) General Purpose -
Huang et al. [41] Cent Homo T Sim General Purpose GPT-4 Turbo
Kannan et al. [50] Cent Homo T Sim(AI2THOR), Real Household GPT-4, GPT-3.5, Llama2, Claude3
Wang et al. [120] Cent Hetero T, I Sim Construction L: Llama-3.1, Claude 3.5, GPT-4o,GPT-3.5 Turbo, V: CLIP
Xu et al. [127] Cent Homo T Sim(AI2THOR+ALFRED), Real Household Minstral, GPT-4
Obata et al. [90] Cent Hetero T Sim(ROS) General Purpose GPT
Liu et al. [74] Hier Hetero T Sim(BEHAVIOR-1K), Real General Purpose GPT-4
Liang et al. [68] Cent Homo T Sim(RoCoBench) General Purpose Llama3.1, Nemotron LLM
Mandi et al. [84] Dec Homo T Sim, Real Household GPT-4
Peng et al. [93] Cent Homo T Sim General Purpose DeepSeek-R1-Distill-Qwen-32B, SFT-Qwen2.5-Coder-7B-Instruct
Yu et al. [135] Dec Hetero T Sim(PyBullet) Household GPT-3.5 Turbo, GPT-4o, Llama3.1
Sueoka et al. [108] Dec Hetero T, I Real Construction V: GPT-4v, L: GPT-4
Hunt et al. [43] Cent Homo T Real General Purpose GPT-4
Yoshida et al. [133] Dec Hetero T, I - General Purpose -
Wang et al. [118] Cent Hetero T Sim(AI2THOR) General Purpose GPT-3.5, Llama-2, Llama-3
Guzman-Merino et al. [38] Cent Hetero T - General Purpose GPT-4o
Cladera et al. [19] Dec Hetero T, I Sim, Real Target Tracking GPT-4o
Chen et al. [14] Cent Homo T Sim Formation GPT-4o, GPT, DeepSeek-R1
Gupta et al. [36] Cent Hetero T Sim Field GPT-4, Gemini, Llama
Wan et al. [115] Cent Hetero T, I Sim, Real Field GPT-4, fine-tuned MultiPlan
Huang et al. [42] Cent Homo T Sim, Real Formation GPT-4.1
Mid level Motion Planning
Yu et al. [134] Cent Homo T Sim(HM3D) Household GPT-3.5 Turbo
Morad et al. [88] Dec Homo T Real General Purpose GTE-Based
Godfrey et al. [31] Cent Homo T Sim, Real General Purpose Llama-3.1
Garg et al. [29] Cent Homo T, I Sim, Real General Purpose V: Claude-3 Sonnet/Opus, GPT-4 Turbo, GPT-4o, L: GPT-4, GPT-3.5, Claude-2, Claude-3
Mahadevan et al. [83] Dec Homo T Sim Game GPT-4o-mini
Wu et al. [124] Cent Homo T Sim(ROS, RViz), Real Target Tracking GPT-4o, GPT-3.5 Turbo
Wu et al. [123] Dec Homo T Sim Household GPT-4o
Shen et al. [104] Dec Homo T, I Sim(HM3D, MP3D) Target Tracking GLM-4V-9B
Ji et al. [46] Dec Homo T Sim Warehouse GPT-3.5, GPT-4
Li and Zhou [64] Dec Homo T Sim, Real Formation o3-mini, Llama3.2, Qwen-Max, DeepSeek-R1, Claude-3.5 Sonnet
Rajvanshi et al. [96] Dec Hetero T Sim Household GPT-4
Wang et al. [119] Dec Homo T Sim Household GPT-4o
Low level Action Execution
Chen et al. [17] Cent Homo T Sim General Purpose GPT-4
Venkatesh et al. [113] Cent Homo T, I Sim(Pygame), Real General Purpose GPT-4, Llama-2, Claude-3 Opus
Li et al. [61] Dec Homo T Sim General Purpose GPT-3.5 Turbo
Strobel et al. [107] Dec Homo T Sim(ARGoS), Real General Purpose GPT-3.5, GPT-4
Lykov et al. [80] Cent Homo T Sim Formation GPT-4
Xue et al. [129] Cent Homo T Sim Formation GPT-4 Turbo
Ji et al. [47] Cent Homo T, I Sim(ROS) Formation o1-mini, GPT-4o, MetaGPT, CaP
Li et al. [66] Dec Hetero T, I Sim Field GPT-4
@@ 5 Application
Others Including Construction, Formation, Target Tracking, and Game. Several studies focused on applications in open-world environments, highlighting the versatility and innovative potential of LLM-integrated robotic systems. For instance, Wang et al. [120] and Sueoka et al. [108] explored the use of LLMs in orchestrating robotic systems for excavation and transportation tasks, showcasing their applicability in construction and complex terrain rescue operations. In drone formations applications, Lykov et al. [80] emphasized the coordination and adaptability needed for outdoor tasks such as search-and-rescue missions and environmental monitoring. Similarly, Wu et al. [124] addressed open-world target tracking by integrating danger zone recognition, providing robust solutions for autonomous navigation in hazardous environments. These scenarios further demonstrate the potential of LLMs in dynamic and structured environments. Brienza et al. [10] introduced LLCoach, a framework for robotic soccer applications, where LLMs enhance strategic decision-making and team coordination. Collectively, these studies underscore the potential of LLM-driven MRS to tackle diverse and complex challenges across various domains.
@@ 6 LLMs, Simulations, and Benchmarks
6.1 LLMs, VLMs, and VLAs
6.1.1 Extending LLM Capabilities
As discussed in Sec. 3-5, LLMs form the backbone of many recent advances in MRS, enabling robots to follow natural language instructions, perform multi-step reasoning, allocate tasks, plan motions, and interact with human operators. Their strengths include flexible problem-solving across diverse scenarios, integration of symbolic or structured data, and the ability to generalize to new missions without retraining.
However, LLMs primarily operate on textual or symbolic representations of the environment. They rely on external perception modules to provide structured inputs such as maps, object lists, or state summaries. This separation between perception and reasoning can limit adaptability in visually complex or fast-changing environments where world states must be updated and integrated in real time.
To address these limitations, researchers have introduced multi-modal extensions, most notably VLMs and Vision-Language-Action models (VLAs) [141, 91]. VLMs incorporate visual information directly into the reasoning process, allowing robots to ground their decisions in raw sensory data [52, 11]. VLAs take this one step further by integrating perception, reasoning, and action generation into a single pipeline, enabling faster, more reactive responses in dynamic settings [11].
6.1.2 VLMs vs. Traditional CV
Traditional computer vision (CV) in MRS is typically task-specific, using separate modules for detection, segmentation, mapping, or action recognition. These components produce structured outputs that are passed to independent reasoning or control systems. While this modularity can yield strong performance for narrow tasks, it often lacks the flexibility and adaptability needed for unstructured or dynamic missions.
VLMs bridge this gap by embedding visual features into a shared semantic space with language, enabling unified perception and reasoning. This allows recognition of previously unseen objects without retraining, switching between perception tasks by altering prompts, and direct connection between sensory input and high-level reasoning without intermediate, hand-coded logic. In MRS, these capabilities have been leveraged for perception-informed coordination, such as in Brienza et al. [10], where a VLM coach transformed soccer match video into schematic scene descriptions for strategic planning, and in the VADER framework [2], where PaLI and CLIP grounded instructions in real-time affordance detection for replanning and recovery.
6.1.3 VLA Models in MRS
VLAs extend VLMs by coupling perception and reasoning directly with action generation [141]. Instead of outputting textual task plans, they can map visual observations and natural language commands to executable control sequences. This closed-loop capability shortens the cycle between sensing and actuation, allowing robots to respond faster to dynamic changes.
In an MRS setting, each VLA-equipped robot could combine the onboard camera feed, interpret the cooperative context, and directly issue motor commands for joint manipulation or navigation without routing through a separate planning layer. Although still relatively new in MRS research, VLAs are being explored in embodied AI for scenarios requiring fast perception–action loops, such as swarm coordination or hazard avoidance.
6.1.4 Comparative Capabilities and Integration Patterns
From an MRS perspective, LLMs, VLMs, and VLAs have complementary strengths. LLMs excel at long-horizon reasoning, symbolic planning, and high-level negotiation between agents. VLMs add environmental grounding through perception, enabling multi-robot teams to share and reason about visual context. VLAs integrate perception and action for rapid, local decision-making.
Integration patterns in the literature generally follow three approaches: (1) visual grounding with language reasoning, where a vision encoder produces scene features that an LLM uses for planning and coordination [10, 29]; (2) direct multi-modal prompting, where visual and textual inputs are processed jointly for end-to-end reasoning [29, 108]; and (3) perception-to-action mapping, where VLA-style architectures produce executable actions directly from multi-modal inputs, bypassing explicit textual reasoning.
Hybrid systems can combine these strengths. For example, a central LLM might manage mission-level strategy, VLMs on individual robots could process onboard sensor data for situational awareness, and VLAs could execute low-latency control in fast-changing conditions.